solve
ySQUARED=13y-42

Answer by checkley75(3666)

Y^2=13Y-42
Y^2-13Y+42=0
(Y-6)(Y-7)=0
Y-6=0
Y=6 ANSWER.
Y-7=0
Y=7 ANSWER.